OpenStudy (anonymous):

lim(3x-15/(sqrt(x)-sqrt(5))) x->5

OpenStudy (anonymous):

Are you allowed L'Hospital? It makes it easy ...

OpenStudy (anonymous):

Its homework, so yeah.

OpenStudy (anonymous):

=lim 3(x-5).(sqrt(x)+sqrt(5))/(x - 5) ... just rationalizing the fraction x->5 =lim 3.(sqrt(x)+sqrt(5)) x->5 = 3 (2sqrt(5) =6sqrt(5)

OpenStudy (anonymous):

I think this is sufficient right?????

OpenStudy (anonymous):

hmm okay lets avoid l'hospital lets do algebra \[\lim_{x\to5} \frac{3(x - 5)}{\sqrt x -\sqrt x }\implies \frac{3(\cancel{(\sqrt x - \sqrt 5)}(\sqrt 5 + \sqrt x)}{\cancel{\sqrt x -\sqrt 5}}\]

OpenStudy (anonymous):

Yep. That's the way.

OpenStudy (anonymous):

\[3( \sqrt 5 + \sqrt 5 ) \implies 3*2\sqrt 5\]

OpenStudy (anonymous):

If you have any doubts then do ask them :]

OpenStudy (anonymous):

It's right. Thanks! :)